A fixed mass of gas is heated inside a rigid container. As its temperature changes from T1


to T2 the pressure increases from 1x105 Pa to 2x105 Pa. Which row in the table shows


possible values for T1 and T2?



T1 T2


A 27°C 327°C


B 30°C 60°C


C 80°C 40°C


D 303 K 333 K


E 600 K 300 K

1
Expert's answer
2021-12-03T12:43:37-0500

First, apply the combined gas law:


"\frac{P_1V}{T_1}=\frac{P_2V}{T_2},\\\space\\
\frac{T_1}{T_2}=\frac{P_1}{P_2}=0.5."

So, the ratio of temperatures in Kelvins should be close to 0.5 as well:


"\frac{27+273}{327+273}=0.5."

The temperatures are 27°C and 327°C.